# MVMS-RCN: A Dual-Domain Unfolding CT Reconstruction with Multi-sparse-view and Multi-scale Refinement-correction

### v1

@@ Abstract
X-ray Computed Tomography (CT) is one of the most important diagnostic imaging techniques in clinical applications. Sparse-view CT imaging reduces the number of projection views to a lower radiation dose and alleviates the potential risk of radiation exposure. Most existing deep learning (DL) and deep unfolding sparse-view CT reconstruction methods: 1) do not fully use the projection data; 2) do not always link their architecture designs to a mathematical theory; 3) do not flexibly deal with multi-sparse-view reconstruction assignments. This paper aims to use mathematical ideas and design optimal DL imaging algorithms for sparse-view tomography reconstructions. We propose a novel dual-domain deep unfolding unified framework that offers a great deal of flexibility for multi-sparse-view CT reconstruction with different sampling views through a single model. This framework combines the theoretical advantages of model-based methods with the superior reconstruction performance of DL-based methods, resulting in the expected generalizability of DL. We propose a refinement module that utilizes unfolding projection domain to refine full-sparse-view projection errors, as well as an image domain correction module that distills multi-scale geometric error corrections to reconstruct sparse-view CT. This provides us with a new way to explore the potential of projection information and a new perspective on designing network architectures. All parameters of our proposed framework are learnable end to end, and our method possesses the potential to be applied to plug-and-play reconstruction. Extensive experiments demonstrate that our framework is superior to other existing state-of-the-art methods. Our source codes are available at https://github.com/fanxiaohong/MVMS-RCN.
@@
Deep learning, deep unfolding network, multi-scale geometric correction, multi-view projection, sparse-view CT reconstruction
@@ I Introduction
where $\bm{y}_{s}\in\mathbb{R}^{q}~{}(q=q_{1}\times q_{2})$ is a sparse-view projection observation, $q_{1}$ and $q_{2}$ are the numbers of views and detector elements, respectively. $\bm{x}\in\mathbb{R}^{d}~{}(d=mn)$ is an unknown image that will be reconstructed. $\bm{P}_{s}\in\mathbb{R}^{q\times d}$ is a sparse-view projection transformation that models the CT imaging system. $\mathcal{W}(\bm{x})$ is a regularizer that incorporates image sparsity prior, and $\lambda$ is a regularization parameter.
@@
The above methods for CT image reconstruction do not verify the consistency of the data with the projection measurement. Several joint dual-domain reconstruction models have been proposed for low-dose CT imaging [28, 29]. CD-Net, which combines a projection domain network, an analytical reconstruction operator, and an image domain network, is designed to restore accurate anatomical information from both the projection and the image domains [30]. DRONE, which integrates priors driven by data in the data and image domains with the kernel awareness used for compressed sensing, has been shown to produce better CT reconstruction results [31]. DL techniques can reduce computational costs and achieve successful CT reconstruction results compared to model-based methods. However, these DL-based methods, which are trained with large amounts of data, are known as an unknown black box with limited theoretical understanding.
@@
There are various ways to formulate and design a convolutional neural network (CNN) model, taking inspiration from CT imaging systems and reconstructing algorithms. LEARN (Learned Experts’ Assessment-Based Reconstruction Network) [32] is a ”fields of experts”-based iterative reconstruction scheme that trains some stages for sparse-projection CT imaging. PD-Net [33] incorporates a forward operator (possibly nonlinear) into a CNN architecture by unrolling a proximal-based primal-dual optimization method, with proximal operators replaced by CNN modules. ISTA-Net and ISTA-Net+ [34] are based on the Iterative Shrinkage-Thresholding Algorithm (ISTA) [35, 36], and learn the proximal mapping associated with the sparsity-inducing regularizer using nonlinear transforms. FISTA-Net [37] unfolds the Fast Iterative Shrinkage/Thresholding Algorithm (FISTA) [38] into a deep network, consisting of multiple gradient descent, proximal mapping (a proximal operator network for nonlinear thresholding), and momentum modules in cascade. AMP-Net [39] is established by unfolding the iterative denoising process of the approximate message passing (AMP) algorithm [40] rather than learning regularization terms. Nest-DGIL [41] based on the second Nesterov proximal gradient optimization has a powerful learning ability for high-/low-frequency image features and can theoretically guarantee the reconstruction of more geometric texture details. These networks combine the advantages of model prior-based and learning-based approaches, such as the interpretability and generality of model-based methods, and the high efficiency and parameter tuning-free advantages of DL-based methods [42, 43, 44, 45, 46, 47, 48, 49].
@@
However, there are some limitations of these imaging methods. 1) the existing methods generally only correct the projection error in single projection view; 2) the derivation from mathematical theory to network design is not always natural enough for these existing deep unfolding methods. The forward and pseudoinverse operators are not well analyzed and explained in PD-Net, just replaced by several convolution layers, also for the CNN denoiser in AMP-Net; 3) there is often a contradiction between real multi-sparse-view scenarios and existing sparse-view CT imaging methods that require re-training for each given scenario.
@@
To overcome these drawbacks, we design a refinement-correction architecture that consists of multi-view projection refinement in the projection domain and multi-scale geometric correction in the image domain. Specifically, 1) we refine projection errors from both the projected and non-projected views in both sparse-view and full-view settings, instead of solely focusing on the error measure from the projected views, giving a new way to explore the potential of available projection information in the projection domain; 2) we transform the completion of the details in the image domain into correcting different scales of errors about the target image geometric errors under multi-scale and multi-resolution. The derivation from mathematical theory to network design is more natural than the existing deep unfolding methods; 3) the proposed unified dual-domain framework can train/predict multi-sparse-view CT reconstruction tasks through a single model to avoid expensive training costs and ample storage space, while the existing sparse-view CT methods have to train an independent model for each sparse-view CT task. The main contributions of this work can be summarized as follows.
@@
    We propose a novel dual-domain deep unfolding unified framework that provides significant flexibility for multi-sparse-view CT reconstruction across different sampling views using a single model. This framework integrates the theoretical advantages of model-based methods with the superior reconstruction performance of DL-based methods, resulting in the expected generalizability of DL.
@@
    We propose a refinement module that utilizes the unfolding projection domain to refine the full-sparse-view projection errors, as well as an image domain correction module that distills multi-scale geometric error corrections to reconstruct the sparse-view CT. This provides us with a new way to explore the potential of projection information and a new perspective on designing network architectures.
@@
    All parameters of our proposed framework are learnable end to end, and our method possesses the potential to be applied to plug-and-play reconstruction. Extensive experiments demonstrate that our framework is superior to other existing state-of-the-art methods.
@@ II Background
where $\sigma_{p}$ is the most commonly used sparse-view projection error extractor.
@@ III-B Multi-scale Geometric Correction Module $\mathcal{D}$
where $\mathcal{E}\left(\mathcal{O}^{n}\right)(\cdot)$ computes the truncation remainder of the operator decomposition. $\mathcal{O}^{i}(\cdot)$ is used to extract different scale geometric errors to correct the intermediate reconstruction $\bm{r}_{\ell}$. Combining different correction operators $\mathcal{O}^{i}$ realizes an optimal regularization prior (such as $\text{Prox}^{\lambda}_{\mathcal{W}}$ in (5)).
@@
where we set $\mathcal{N}_{n}=\mathcal{G}_{n+1}$ in this study. Many deep unfolding networks that take a single channel image as input have been shown to significantly hinder information transmission and lose many image details [54]. So we first transform the intermediate reconstruction $\bm{r}_{\ell}$ into features with $p$ channels through the feature extractor $\mathcal{C}_{m}\left(\cdot\right)$, which consists of convolution (with $p$ channels and a 3$\times$3 kernel) and LeakyReLU. Furthermore, the feature fusion operator $\mathcal{C}_{a}\left(\cdot\right)$, which consists of convolution (with $p$ channels and a 3$\times$3 kernel) and LeakyReLU, is used to recover corrected high-throughput information in a reconstructed CT image.
@@
1:$\mathcal{L}_{\text{total}}(\bm{\Theta})=\mathcal{L}_{\text{pixel}}(\bm{\Theta})+\gamma\mathcal{L}_{\text{SSIM}}(\bm{\Theta});$
@@ III-C Loss Function
Sparse-view projection data $\bm{y}_{s}$, downsampling factor $\alpha$ and initialization $\bm{x}_{0}$ are used as inputs to obtain the final reconstructed output $\bm{x}_{n_{s}}$ through the proposed method. The loss function is then used to find the target image $\bm{x}^{*}$ by minimizing the distance between $\bm{x}_{n_{s}}$ and the original image without artifacts $\bm{x}_{f}$. Here, we employ a $\ell_{1}$ loss instead of a $\ell_{2}$ loss which is not sufficient to capture perceptually relevant components (e.g., high-frequency geometric details) [55], to increase the original pixel loss. Furthermore, we use a Structural Similarity Index Measure (SSIM) loss to quantify differences in brightness, contrast, and structure between $\bm{x}_{n_{s}}$ and $\bm{x}_{f}$, to help train the network. The total loss is expressed as follows:
$\begin{split}\mathcal{L}_{\text{total}}(\bm{\Theta})&=\mathcal{L}_{\text{pixel}}(\bm{\Theta})+\gamma\mathcal{L}_{\text{SSIM}}(\bm{\Theta})\\ &=\frac{1}{mn}\left\|\bm{x}_{\scriptscriptstyle n_{s}}-\bm{x}_{f}\right\|_{1}+\gamma(1-{\text{SSIM}}\left(\bm{x}_{\scriptscriptstyle n_{s}},\bm{x}_{f}\right)),\end{split}$
where $\gamma=1$ is the weight coefficient. $mn$ is the size of $\bm{x}_{f}$, and $n_{s}$ is the total stage number of the proposed method.
@@ IV-A Implementation Details
For fan-beam projection, we employ a two-dimensional fan-beam geometry with 1024 angles evenly distributed over a full 360-degree radius. The detector is a plane that contains 1024 pixels that are contiguous and spaced 2mm apart. The projection data with $1024\times 1024$ resolutions are generated by full-dose CT sampling using the RadonFanbeam operator in TorchRadon [58]. These data can be down-sampled to 16, 32, 64, 128 and 256 views to simulate sparse-view geometries. However, the original artifact-free images are reconstructed by FBP using all 1024 views (full view).
@@ IV-B Intra-Method Evaluation
We perform five groups of experiments to determine the configuration of the proposed MVMS-RCN for sparse-view CT reconstruction. These experiments involve the number of stages $n_{s}$, the multiple scales depth $n$ in each stage, different initial inputs, an ablation study and the sharing of modules.
@@ IV-B1 Stage number $n_{s}$
We assess the reconstruction performance of different stage numbers by varying the stage number $n_{s}$ from 5 to 9. Table I shows the mean PSNR and SSIM for the fan-beam projection dataset with different views when the frameworks are used with different stage numbers. To balance the complexity of the framework and the reconstruction performance [37], we fix $n_{s}=7$ for all experiments.
@@ IV-B2 Multi-scale depth $n$
We investigate the effect of the multi-scale geometric error correction depth $n$ on the reconstruction performance. We vary the depth $n$ from 2 to 6. The mean PSNR and SSIM values of the reconstructed CT images at different depths are summarized in Table I. It can be observed that the reconstruction performance gradually increases as the depth increases, up to $n=5$. After this point, the performance becomes stable.We selected a depth of 5 for all configurations, as it provides a good balance between network complexity and reconstruction performance.
@@ IV-B3 Different initial inputs
In this part, we evaluate the influence of two initialization schemes, $\bm{x}_{0}=\bm{P}_{s}^{T}\bm{y}_{s}$ and $\bm{x}_{0}=0$, on the proposed MVMS-RCN (fan-beam projection with 32 views). Table II shows the quantitative assessment of the different initial inputs $\bm{x}_{0}$. The reconstruction performance of $\bm{x}_{0}=0$ is slightly less than that of $\bm{x}_{0}=\bm{P}_{s}^{T}\bm{y}_{s}$, which validates the suitability of the initialization technique we employ, akin to numerous existing methods. Nevertheless, both initialization schemes yield good reconstruction performances, due to the effective error extraction of multi-view projection refinement and multi-scale geometric correction.
TABLE III: Quantitative assessment with PSNR values of different combinations of projection errors by proposed MVMS-RCN.
@@
IV-B4 Ablation study
Next, we conduct a group of ablation studies to evaluate the effectiveness of the SVR error $(\bm{e}_{\ell,s})$, SVP errors $(\bm{e}_{\ell,d},\bm{e}_{\ell,j})$, FVP errors $(\bm{e}_{\ell,f},\bm{e}_{\ell,k})$ and CVR $(\bm{x}_{u},\bm{e}_{\ell,u})$ on the performance of the CT reconstruction with sparse view. Comparisons are shown in Table III. Unlike variant (a) which only relies on the previous output $\bm{x}_{\ell-1}$, variant (b) can use CVR reconstruction and error $(\bm{x}_{u},\bm{e}_{\ell,u})$ to greatly enhance CT reconstruction performance. From the comparison between variants (a) and (e) or the comparison between variants (c) and (d), the effectiveness of the widely used SVR error is apparent. From the comparison between variants (e) and (f) or the comparison between variants (d) and (g), the proposed SVP errors can further enhance the CT reconstruction performance. The comparison of variants (b) and (c) reveals that the proposed FVP errors can effectively improve CT reconstruction performance. From the above observations, we find that the proposed full-sparse-view projection error refinement techniques can effectively extract the potential projection information.
TABLE IV: Quantitative assessment with different shared settings of MVMS-RCN.

### v2

# MVMS-RCN: A Dual-Domain Unified CT Reconstruction with Multi-sparse-view and Multi-scale Refinement-correction

@@ Abstract
X-ray Computed Tomography (CT) is one of the most important diagnostic imaging techniques in clinical applications. Sparse-view CT imaging reduces the number of projection views to a lower radiation dose and alleviates the potential risk of radiation exposure. Most existing deep learning (DL) and deep unfolding sparse-view CT reconstruction methods: 1) do not fully use the projection data; 2) do not always link their architecture designs to a mathematical theory; 3) do not flexibly deal with multi-sparse-view reconstruction assignments. This paper aims to use mathematical ideas and design optimal DL imaging algorithms for sparse-view CT reconstructions. We propose a novel dual-domain unified framework that offers a great deal of flexibility for multi-sparse-view CT reconstruction through a single model. This framework combines the theoretical advantages of model-based methods with the superior reconstruction performance of DL-based methods, resulting in the expected generalizability of DL. We propose a refinement module that utilizes unfolding projection domain to refine full-sparse-view projection errors, as well as an image domain correction module that distills multi-scale geometric error corrections to reconstruct sparse-view CT. This provides us with a new way to explore the potential of projection information and a new perspective on designing network architectures. The multi-scale geometric correction module is end-to-end learnable, and our method could function as a plug-and-play reconstruction technique, adaptable to various applications. Extensive experiments demonstrate that our framework is superior to other existing state-of-the-art methods. Our source codes are available at https://github.com/fanxiaohong/MVMS-RCN.
@@
Deep learning, unfolding explainable network, multi-scale geometric correction, multi-view projection, sparse-view CT reconstruction, plug-and-play.
@@ I Introduction
where $\bm{y}_{s}\in\mathbb{R}^{q}~{}(q=q_{1}\times q_{2})$ is a sparse-view projection observation, $q_{1}$ and $q_{2}$ are the numbers of views and detector elements, respectively. $\bm{x}\in\mathbb{R}^{d}$ is an unknown image to be reconstructed, where $d=m_{1}\times m_{2}$, and $m_{1}$ and $m_{2}$ are the numbers of rows and columns of $x$, respectively. $\bm{P}_{s}\in\mathbb{R}^{q\times d}$ is a sparse-view projection transformation that models the CT imaging system. $\mathcal{W}(\bm{x})$ is a regularizer that incorporates image sparsity prior. $\lambda$ is a regularization parameter that should be determined based on the specific scenario.
@@
The aforementioned CT image reconstruction methods do not utilize the prior information from the known projection observation $\bm{y}_{s}$, which limits their reconstruction performance. Several joint dual-domain reconstruction models have been proposed for low-dose CT imaging [28, 29]. CD-Net, which combines a projection domain network, an analytical reconstruction operator, and an image domain network, is designed to restore accurate anatomical information from both the projection and the image domains [30]. DRONE, which integrates data-driven priors with kernel awareness used for compressed sensing, has been shown to produce better CT reconstruction results [31]. DL techniques can reduce computational costs and achieve successful CT reconstruction results compared to model-based methods. However, these DL-based methods, which are trained with large amounts of data, are known as an unknown black box with limited theoretical understanding.
@@
There are various ways to formulate and design a convolutional neural network (CNN) model, taking inspiration from CT imaging systems and reconstructing algorithms. LEARN (Learned Experts’ Assessment-Based Reconstruction Network) [32] is a ”fields of experts”-based iterative reconstruction scheme that trains some stages for sparse-projection CT imaging. PD-Net [33] incorporates a forward operator (possibly nonlinear) into a CNN architecture by unrolling a proximal-based primal-dual optimization method, with proximal operators replaced by CNN modules. ISTA-Net and ISTA-Net+ [34] are based on the Iterative Shrinkage-Thresholding Algorithm (ISTA) [35, 36], and learn the proximal mapping associated with the sparsity-inducing regularizer using nonlinear transforms. FISTA-Net [37] unfolds the Fast Iterative Shrinkage/Thresholding Algorithm (FISTA) [38] into a deep network, consisting of multiple gradient descent, proximal mapping (a proximal operator network for nonlinear thresholding), and momentum modules in cascade. AMP-Net [39] is established by unfolding the iterative denoising process of the approximate message passing (AMP) algorithm [40] rather than learning regularization terms. Nest-DGIL [41] based on the second Nesterov proximal gradient optimization has a powerful learning ability for high-/low-frequency image features and can theoretically guarantee the reconstruction of more geometric texture details. DIVA [42] unfolded a baseline adaptive denoising algorithm (DeQuIP[43, 44]), relying on the theory of quantum many-body physics. It can handle non-local image structures through the patch-interaction term and the quantum-based Hamiltonian operator. DPIR [45] plugs a deep denoiser prior into a half quadratic splitting based iterative algorithm to solve various image restoration problems. These networks combine the advantages of model prior-based and learning-based approaches, such as the interpretability and generality of model-based methods, and the high efficiency and parameter tuning-free advantages of DL-based methods [46, 47, 48, 49, 50, 51, 52, 53].
@@
However, there are some limitations of these imaging methods. 1) the existing methods generally only correct the projection error in single projection view; 2) the derivation from mathematical theory to network design is not always natural enough for these existing deep unfolding methods. In PD-Net, the forward and pseudoinverse operators lack thorough analysis and explanation, being substituted by CNNs. This problem is similarly found in the CNN denoisers utilized in AMP-Net and DPIR, and in the inversion of the degradation operator in DIVA. These mathematical operations are replaced by CNN without a clear theoretical foundation; 3) there is often a contradiction between real multi-sparse-view scenarios and existing sparse-view CT imaging methods that requires re-training for each given scenario.
@@
To overcome these drawbacks, we design a refinement-correction architecture that consists of multi-view projection refinement in the projection domain and multi-scale geometric correction in the image domain. Specifically, 1) we refine projection errors from both the projected and non-projected views in both sparse-view and full-view settings, instead of solely focusing on the error measure from the projected views, giving a new way to explore the potential of available projection information in the projection domain; 2) we transform the completion of the details in the image domain into correcting different scales of errors about the target image geometric errors under multi-scale and multi-resolution. By aligning the detailed mathematical framework with the network architecture, our approach ensures that all mathematical operations are systematically integrated into the network design. The derivation from mathematical theory to network design is more natural than the existing deep unfolding methods; 3) the proposed unified dual-domain framework can train/predict multi-sparse-view CT reconstruction tasks through a single model to avoid expensive training costs and ample storage space, while the existing sparse-view CT methods have to train an independent model for each sparse-view CT task. The main contributions of this work can be summarized as follows.
@@
    We propose a novel dual-domain unified framework that provides significant flexibility for multi-sparse-view CT reconstruction through a single model. This framework integrates the theoretical advantages of model-based methods with the superior reconstruction performance of DL-based methods, resulting in the expected generalizability of DL.
@@
    We propose a multi-view projection refinement module $\mathcal{R}$ that utilizes the unfolding projection domain to refine the full-sparse-view projection errors to reconstruct the multi-sparse-view CT. This provides us with a new way to explore the potential of projection information.
@@
    We propose an multi-scale geometric correction module $\mathcal{D}$ that distills multi-scale geometric error corrections to reconstruct the multi-sparse-view CT. This provides us with a new perspective on designing network architectures that systematically integrate mathematical operations.
  • (4)
    The multi-scale geometric correction module $\mathcal{D}$ is learnable end to end, and our method could serve as a plug-and-play reconstruction technique suitable for various applications. Extensive experiments demonstrate that our framework is superior to other existing state-of-the-art methods.
@@ II Background
where $\sigma_{p}$ is the most commonly used operator for consistency checking of projection data.
@@ III-B Multi-scale Geometric Correction Module $\mathcal{D}$
where $\mathcal{E}\left(\mathcal{O}^{n}\right)(\cdot)$ computes the truncation remainder of the operator decomposition. $\mathcal{O}^{i}(\cdot)$ is used to extract different scale geometric errors to correct the intermediate reconstruction $\bm{r}_{\ell}$. $n$ is the multi-scale geometric error correction depth. In this study, we realize an optimal regularization prior (such as $\text{Prox}^{\lambda}_{\mathcal{W}}$ in (5)) by combining different correction operators $\mathcal{O}^{i}$, thereby removing the necessity to determine a fixed $\lambda$.
@@
where $\mathcal{C}_{m}\left(\cdot\right)=\operatorname{LeakyReLU}\left(\operatorname{Conv}\left[\cdot;1,p\right]\right)$ is a feature extractor. $\mathcal{C}_{a}\left(\cdot\right)=\operatorname{LeakyReLU}\left(\operatorname{Conv}\left[\cdot;p,1\right]\right)$ is a feature fusion operator. $\operatorname{Conv}\left[\cdot;p^{\mathrm{in}},p^{\text{out }}\right]$ is the $3\times 3$ convolution. We set $\mathcal{N}_{n}=\mathcal{G}_{n+1}$ in this study. Many deep unfolding networks that take a single channel image as input have been shown to significantly hinder information transmission and lose many image details [58]. So we first transform the intermediate reconstruction $\bm{r}_{\ell}$ into features with $p$ channels through $\mathcal{C}_{m}\left(\cdot\right)$. Then we recover the corrected high-throughput information to a reconstructed CT image by $\mathcal{C}_{a}\left(\cdot\right)$.
@@
1:$\mathcal{L}_{\text{total}}(\bm{\Theta})=\mathcal{L}_{\text{pixel}}(\bm{\Theta})+\gamma\mathcal{L}_{\text{SSIM}}(\bm{\Theta});$ (13)
@@ III-C Loss Function
Sparse-view projection data $\bm{y}_{s}$ and initialization $\bm{x}_{0}$ are used as inputs to obtain the final reconstructed output $\bm{x}_{n_{s}}$ through the proposed method. The loss function is then used to find the target image $\bm{x}^{*}$ by minimizing the distance between $\bm{x}_{n_{s}}$ and the original image without artifacts $\bm{x}_{f}$. Here, we employ a $\ell_{1}$ loss instead of a $\ell_{2}$ loss which is not sufficient to capture perceptually relevant components (e.g., high-frequency geometric details) [59], to increase the original pixel loss. Furthermore, we use a Structural Similarity Index Measure (SSIM) loss to quantify differences in brightness, contrast, and structure between $\bm{x}_{n_{s}}$ and $\bm{x}_{f}$, to help train the network. The total loss is expressed as follows:
$\mathcal{L}_{\text{total}}(\bm{\Theta})=\mathcal{L}_{\text{pixel}}(\bm{\Theta})+\gamma\mathcal{L}_{\text{SSIM}}(\bm{\Theta})$ (13)
$\text{with}:\begin{cases}\mathcal{L}_{\text{pixel}}(\bm{\Theta})=\frac{1}{d}\left\|\bm{x}_{\scriptscriptstyle n_{s}}-\bm{x}_{f}\right\|_{1}\\ \mathcal{L}_{\text{SSIM}}(\bm{\Theta})=1-{\text{SSIM}}\left(\bm{x}_{\scriptscriptstyle n_{s}},\bm{x}_{f}\right),\end{cases}$
where $n_{s}$ is the total stage number of the proposed method. We set the weight coefficient $\gamma=1$.
@@ IV-A Implementation Details
For fan-beam projection, we employ a two-dimensional fan-beam geometry with 1024 angles evenly distributed over a full 360-degree radius. The detector is a plane containing 1024 pixels that are contiguous and spaced 2 mm apart. Both the source and the detector are positioned 500 mm away from the rotation center. The projection data with $1024\times 1024$ resolutions are generated by full-dose CT sampling using the RadonFanbeam operator in TorchRadon [62]. These data can be down-sampled to 16, 32, 64, 128 and 256 views to simulate sparse-view geometries. The original artifact-free images are reconstructed by FBP using all 1024 views (full view).
@@ IV-B Intra-Method Evaluation
We perform several groups of experiments to determine the optimal configuration of the proposed MVMS-RCN for sparse-view CT reconstruction. These experiments involve the number of stages $n_{s}$, the depth of multiple scales $n$ within each stage, different initial inputs, different loss weight, an ablation study, and the sharing of modules.
@@ IV-B1 Stage number $n_{s}$
We assess the reconstruction performance of different stage numbers by varying the stage number $n_{s}$ from 5 to 9. Table I shows the average PSNR for the fan-beam projection dataset with different views when the proposed framework is implemented with different stage numbers. To balance the complexity of the framework and the reconstruction performance [37], we therefore fix $n_{s}=7$ for all subsequent experiments.
@@ IV-B2 Multi-scale depth $n$
We investigate the effect of the multi-scale geometric error correction depth $n$ on the reconstruction performance. We vary the depth $n$ from 2 to 6. Table I presents the average PSNR values for the reconstructed CT images at different depths. It can be observed that the reconstruction performance gradually increases as the depth increases, up to $n=5$. After this point, the performance becomes stable. We selected a depth of $n=5$ for optimal configurations, as it provides a good balance between network complexity and reconstruction performance.
@@ IV-B3 Different initial inputs
In this part, we evaluate the influence of two initialization schemes, $\bm{x}_{0}=\bm{P}_{s}^{T}\bm{y}_{s}$ and $\bm{x}_{0}=0$, on the proposed MVMS-RCN (fan-beam projection with 32 views). Table II shows the quantitative assessment of the different initial inputs $\bm{x}_{0}$. The reconstruction performance of $\bm{x}_{0}=0$ is slightly less than that of $\bm{x}_{0}=\bm{P}_{s}^{T}\bm{y}_{s}$, which validates the suitability of the initialization technique we employ, similar to numerous existing methods. Nevertheless, both initialization schemes yield good reconstruction performances, due to the effective error extraction of multi-view projection refinement and multi-scale geometric correction.
TABLE III: Quantitative assessment with different loss weight coefficient $\gamma$ of MVMS-RCN.
Loss weight Views Avg.
16 32 64 128 256
$\gamma=0$ 38.68 41.71 43.31 44.88 47.37 43.19
$\gamma=0.1$ 38.80 41.74 43.29 44.85 47.31 43.20
$\gamma=1$ 38.98 41.72 43.28 44.85 47.27 43.22
$\gamma=10$ 38.27 41.44 43.24 44.86 47.36 43.03
TABLE IV: Quantitative assessment with PSNR values of different combinations of projection errors by proposed MVMS-RCN.
@@
IV-B4 Different loss weight
The SSIM loss measures luminance, contrast, and structural similarity, which aligns more closely with human visual perception than the $\ell_{1}$ loss. Typically, the SSIM loss results in more detailed reconstructions. In this study, the proposed approach addresses multi-sparse-view CT reconstruction challenges using a single model, making it impractical to fine-tune the weight parameter $\gamma$ for each specific task. Various combinations of weights for these losses were tested, and the results are shown in Table III. Our results suggest that setting $\gamma$ to 1 is optimal.
IV-B5 Ablation study
Next, we conduct a group of ablation studies to evaluate the effectiveness of the SVR error $(\bm{e}_{\ell,s})$, SVP errors $(\bm{e}_{\ell,d},\bm{e}_{\ell,j})$, FVP errors $(\bm{e}_{\ell,f},\bm{e}_{\ell,k})$ and CVR $(\bm{x}_{u},\bm{e}_{\ell,u})$ on the performance of the CT reconstruction with sparse view. Comparisons are shown in Table IV. Unlike variant (a) which only relies on the previous output $\bm{x}_{\ell-1}$, variant (b) can use CVR reconstruction and error $(\bm{x}_{u},\bm{e}_{\ell,u})$ to greatly enhance CT reconstruction performance. From the comparison between variants (a) and (e) or the comparison between variants (c) and (d), the effectiveness of the widely used SVR error is apparent. From the comparison between variants (e) and (f) or the comparison between variants (d) and (g), the proposed SVP errors can further enhance the CT reconstruction performance. The comparison of variants (b) and (c) reveals that the proposed FVP errors can effectively improve CT reconstruction performance. From the above observations, we find that the proposed full-sparse-view projection error refinement techniques can effectively extract the potential projection information.
TABLE V: Quantitative assessment with different shared settings of MVMS-RCN.